OpenStudy (anonymous):

17-9m=n-23, solve for n

OpenStudy (anonymous):

n = 40-9m

OpenStudy (anonymous):

17-9m=n-23 => n = 40-9m if m & n are integers then values of n can be 31,22,13,............... If the m & n are real numbers then there are uncountable values of n